Shift report 20011006 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Oct 06 2001 - 15:50:35 EDT

  • Next message: Apache: "Shift report 20011007 16:00-24:00"

    ------------------------------------------------------------
    Date: Sat Oct  6 15:50:35 2001, Shift: 08:00-16:00
    Supervisor: R. Debbe, with: none
    Log book#9 pages 105 - 106
    
    Shift Summary:
    ------------------------------------------------------------
    Machine status: Store started at 07:25 died at 9:10 by 
    massive QLI At 9:45 MCR announced Beta start squeeze studies
    that are still under way.
    
    Achievements: Run 5193 233K trigg 1
    
    Detector problems: Nothing new, pedestals in TPM2 are back to
    sigmas of 1.5 - 2 channels. Yesterday Eun-Joo took data with 
    DVMs back in the data but DVM1 out of the trigger. Peaks are
    cleanly seen in the other two active DVMs. I will move DVM1
    behind D5 during next access, and then we can resume monitoring
    the drift velocity while we acquire data.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Oct 06 2001 - 15:51:09 EDT